A computerized driving test held ten sessions in a day. The overall passing rate of the driving test is 87.5%.
In each session, there are 14 candidates.
(a) Most likely, how many candidates would pass the test in a session? What is the probability of its
happening?
(b) What is the probability that more than 10 candidates in a session can pass the test?
(c) What are the (i) expectation, (ii) variance, and (iii) standard deviation of number of candidates can pass
the driving test in a day.
(d) Suppose each candidate who fail the driving test will immediately join the tutorial class for the 2nd test.
The charge of the tutorial class is $400. What are the (i) expectation and (ii) standard deviation of the
daily income earned from the tutorial class?
1
Expert's answer
2020-04-03T17:56:17-0400

Let "X=" the number of candidates passing the test: "X\sim Bin(n, p)."


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "p=0.875, n=14"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & P(X=x) \\ \hline
 0 & 2.27\times10^{-13} \\ \hline
 1 & 2.23\times10^{-11} \\ \hline
 2 & 1.01\times10^{-9} \\ \hline
 3 & 2.84\times10^{-8} \\ \hline
 4 & 5.47\times10^{-7} \\ \hline
 5 & 7.65\times10^{-6} \\ \hline
 6 & 8.03\times10^{-5} \\ \hline
 7 & 6.43\times10^{-4} \\ \hline
 8 & 0.003936 \\ \hline
 9 & 0.018369\\ \hline
 10 & 0.064292\\ \hline
 11 & 0.163652\\ \hline
 12 & 0.286390 \\ \hline
 13 & 0.308420 \\ \hline
 14 & 0.154210 \\ \hline
\end{array}"

(a) Most likely, 13 candidates would pass the test in a session. The probability is "0.308420."


(b) The probability that more than 10 candidates in a session can pass the test is


"P(X>10)=P(X=11)+P(X=12)+""+P(X=13)+P(X=14)\approx""\approx0.163652+0.286390+0.308420+0.154210\approx""\approx0.912672"

(c)


"E(X)=np=14(0.875)=125.25"

"Var(X)=\sigma^2=np(1-p)=14(0.875)(1-0.875)=1.53125"

"\sigma=\sqrt{\sigma^2}=\sqrt{1.53125}\approx1.23744"

d)

(i)


"E(Y)=400(14)(1-0.875)=\$700"

(ii)


"Var(Y)=\sigma_Y^2=(400)^2(14)(1-0.875)(0.875)=245000"

"\sigma_Y=\sqrt{245000}=\$494.97"